a flowerbed is in the shape of a rectangle of which one side is 5 m and the other side is 4 m. a rectangular path around the bed has a width of 1,2 m. calculate how many square tiles of side 200 mm will be needed to tile the path
Answer by macston(5194)

.
Outside dimension of path:
Length=path+flower bed+path=1.2m+5m+1.2m=7.4m
Width=path+flower bed+path=1.2m+4m+1.2m=6.4m
.
Area of path and flower bed=(7.4m)(6.4m)=47.36m^2
Area of flower bed=(4m)(5m)=20m^2
Area of path=area of path and flower bed-area of flower bed
Area of path=47.36m^2-20m^2=27.36m^2
.
Area of each tile=(0.200m)(0.200m)=0.04m^2
.
Number of tiles=Area to be tiled/size of tile
Number of tiles=27.36m^2/0.04m^2=684 tiles
.
ANSWER: 684 tiles will be needed.